OpenStudy (anonymous):

4 = 10/2X - 1/X SHOW YOUUR WORK Xd

OpenStudy (anonymous):

10/2x - (1/x)(2/2) (10-2x) / 2x = 4 (10-2x) = 8x 10 = 10x x = 1